OpenStudy (anonymous):

Use the function to answer the question. f(x)=3x; g(x)=4x^2 + 1 Which is equal to (f○g)(x)? a.12x^3+3x b. 12x^2 +3 c.36x^2 +1 d.4x^2 + 3x+1

OpenStudy (dayakar):

f0g(x)= f[g(x)] = f[4x^2+1]

OpenStudy (dayakar):

= 3[4x^2+1]

OpenStudy (anonymous):

so f0g(x)= 3(4x^2+1) = 3(16+1)? = 3(17) = 51?

OpenStudy (dayakar):

is there any x value given

OpenStudy (anonymous):

no.

OpenStudy (dayakar):

so , 12x^2+3

OpenStudy (dayakar):

B is the answer

OpenStudy (dayakar):

is it clear

OpenStudy (anonymous):

because you multply 4 times 3 which gave you 12x^2. then 1 times 3 equal 3. I see now. Thank you

OpenStudy (dayakar):

ok, it's alright